# Anchor-based Multi-view Subspace Clustering with Hierarchical Feature Descent

### v1

@@ 3.2.1 Optimization Algorithm
The decision variables for the objective function are $\{\mathbf{W}^{(v)}_{i}\}_{i=1}^{\delta},\mathbf{A}$, $\mathbf{Z}$ and $\boldsymbol{\alpha}$. As a result, we introduce a four-step iterative optimization method to approximate the optimal target mentioned in Eq. (6). In this approach, we keep three variables fixed while optimizing the remaining ones. The specific procedure is derived from the following detailed steps.
Initialization. Our algorithm adopts a straightforward initialization by setting $\mathbf{Z}$ as an $m\times n$ matrix with the first $m\times m$ block deposited as identity matrix and the rest elements deposited to be 00. The anchor matrix $\mathbf{A}$ is initialized as identity matrix w.r.t. its dimension. And $\boldsymbol{\alpha}$ is initialized using mean values factorized by view count. The other decision variable $\{\mathbf{W}^{(v)}_{i}\}_{i=1}^{\delta}$ is neglected for initializing as it is the first to optimize.
Update $\mathbf{W}^{(v)}_{o}$ with $\{\mathbf{W}^{(v)}_{i}\}_{i=1,i\neq o}^{\delta}$, $\mathbf{A}$, $\mathbf{Z}$ and $\boldsymbol{\alpha}$ fixed. Given $\mathbf{A}$, $\mathbf{Z}$, $\boldsymbol{\alpha}$ and the rest $\delta-1$ numbers of $\{\mathbf{W}^{(v)}_{i}\}_{i=1,i\neq o}^{\delta}$, the optimization problem in Eq. (6) with respect to $\mathbf{W}^{(v)}_{o}$ becomes:
@@
Denote the SVD(Singular Value Decomposition) of matrix $\mathbf{M}^{(v)}$ as $\mathbf{U}\mathbf{D}\mathbf{V}^{\top}$ and the optimal solution of Eq. (9) is demonstrated as ${\mathbf{W}^{(v)}_{o}}^{\star}$. In this case, $\mathbf{D}$ is a nonnegative diagonal matrix with each entry representing the singular value of $\mathbf{M}^{(v)}$, both $\mathbf{U}$ and $\mathbf{V}$ are unitary matrix with orthogonal entries. According to deduction in Ou et al. (2020), a closed-form resolution can be found by calculating ${\mathbf{W}^{(v)}_{o}}^{\star}=\mathbf{U}sign(\mathbf{D})\mathbf{V}^{\top}$, where $sign(\cdot)$ is a function imposed on matrix with each element, which returns $1$ if the element is larger than 00. And it is obvious that in our case, $sign(\mathbf{D})$ is an identity matrix. Accordingly, the optimal solution can be calculated as:
${\mathbf{W}^{(v)}_{o}}^{\star}=\mathbf{U}\mathbf{V}^{\top}$ (10)
By following the updating expression in Eq. (10), our algorithm updates $\mathbf{W}^{(v)}_{i},~{}i=1,2,\cdots,\delta$ in succession.
Update $\mathbf{A}$ with $\{\mathbf{W}^{(v)}_{i}\}_{i=1}^{\delta}$, $\mathbf{Z}$ and $\boldsymbol{\alpha}$ fixed. When $\boldsymbol{\alpha}$ and $\{\mathbf{W}^{(v)}_{i}\}_{i=1}^{\delta}$ of each view are all prescribed, updating $\mathbf{A}$ with fixed $\mathbf{Z}$, which relocates the unified anchor matrix, can be termed as:
@@
where $\Phi=\sum_{v=1}^{p}\alpha_{v}^{2}{\mathbf{W}^{(v)}_{\delta}}^{\top}{\mathbf{W}^{(v)}_{\delta-1}}^{\top}\cdots{\mathbf{W}^{(v)}_{1}}^{\top}\mathbf{X}^{(v)}\mathbf{Z}^{\top}$ and $\Phi\in\mathbb{R}^{k\times m}$.
Corresponding to the updating rule in Ou et al. (2020), it is straightforward to obtain the optimal solution for $\mathbf{A}$ as in Eq. (9). The multiplication of the left and right singular matrices of $\Phi$ costs $\mathcal{O}(km^{2})$.
Update $\mathbf{Z}$ with $\mathbf{A}$, $\{\mathbf{W}^{(v)}_{i}\}_{i=1}^{\delta}$ and $\boldsymbol{\alpha}$ fixed. Given $\{\mathbf{W}^{(v)}_{i}\}_{i=1}^{\delta}$, the anchors are separately lodged in $p$ different anchor spaces of the same dimension $k$ and are reunited by the optimized anchor matrix $\mathbf{A}$, which is shared amongst views. This subtask of updating $\mathbf{Z}$ is equivalent to the process of constructing the bipartite graph from the anchor space to the original feature space. The optimization problem in Eq. (6) concerning $\mathbf{Z}$ can be written as:
@@
for $j=1,2,\cdots,n$, where the parameter for quadratic term is $\mathbf{Q}=2\sum_{v=1}^{p}{\alpha_{v}^{2}}\mathbf{I}$ and the first order polynomial term ${\boldsymbol{q}^{\top}}=-2\sum_{v=1}^{p}{\alpha_{v}^{2}}{\mathbf{X}^{(v)}_{:,j}}^{\top}\mathbf{P}$. These are several QP(quadratic programming) problems with domain constraints. As each column of $\mathbf{Z}$ is an $m$ dimensional vector, and there are $n$ sub-problems identical to each other, the overall time cost of optimising $\mathbf{Z}$ is $\mathcal{O}(nm^{3})$.
Update $\boldsymbol{\alpha}$ with $\mathbf{Z}$, $\mathbf{A}$ and $\{\mathbf{W}^{(v)}_{i}\}_{i=1}^{\delta}$ fixed. When variables $\mathbf{Z}$, $\mathbf{A}$ and $\{\mathbf{W}^{(v)}_{i}\}_{i=1}^{\delta}$ are fixed, the optimization problem in Eq. (6) with regard to the variable $\boldsymbol{\alpha}$ automatically learns the coefficients of different views, the procedure is assembled below:
@@ 4.1.1 A Brief Introduction of Datasets
For in-depth experimental analysis, we conduct a series of tests over 10 real-world datasets, varying in multiple domains. The datasets used in our experiments are WebKB  [^1] http://lig-membres.imag.fr/grimal/data.html, BDGP  [^2] https://www.fruitfly.org/, MSRCV1 Winn and Jojic (2005), Caltech7-2view  [^3] http://www.vision.caltech.edu/Image_Datasets/Caltech101/, NUS-WIDE-10  [^4] https://lms.comp.nus.edu.sg/wp-content/uploads/2019/research/nuswide/NUS-WIDE.html, Caltech101-7 ${}^{3}$, Caltech101-20 ${}^{3}$, NUS-WIDE ${}^{4}$, SUNRGBD  [^5] https://rgbd.cs.princeton.edu/, YoutubeFace  [^6] https://www.cs.tau.ac.il/ wolf/ytfaces/. In Table 2, we summarize the detailed information of these benchmark datasets.
@@ 4.1.2 Experimental Settings
We use the initialization specified in Algorithm 1. In practice, we simply adopt the number of clusters $k$ as the value of the dimension of embedding space and the value of anchor count. For each view, the hierarchical feature descent rate is different according to their original dimensions. However, it is rather practicable as it performs well by directly slicing the dimension to equal sections: $l_{o}-l_{o-1}=l_{o+1}-l_{o},\forall o\in[1,\delta]$.
@@ 4.1.3 Compared Algorithms
RMKM LMVSC AMGL BMVC FMR MVSC SFMC FPMVS-CAG EMKMC
#Parameter 1 2 0 3 2 3 1 0 p
@@
As shown in Table.3, most of algorithm have more than one hyper-parameters. It is encouraged to use learnable parameters which can adaptively collaborate with the optimization of other decision variables to boost the performance whereas spare the acquisition of prior knowledge. For the fast algorithm EMKMC, a parameter is searched recurrently for each view, so the parameter count is equal to its view count.
@@ 4.2 Comparison With State-of-the-Art Algorithms
In this section, we conduct a comparative analysis between our algorithm and several state-of-the-art MVC algorithms to validate its effectiveness. We evaluate the performance of our algorithm using 9 widely-used methods, considering accuracy (ACC), Normalized Mutual Information (NMI), and purity as evaluation metrics. The results are presented in Table 1. It shows that our findings entail most of the best practices and recommendations that can improve multi-view learning performance.
@@ 4.2.1 Running Time Comparison
The running time of compared algorithms on benchmark datasets are shown in Figure 3.
@@
It is worth noticing that for large-scale datasets, namely NUS-WIDE, SUNRGBD and YoutubeFace, some algorithms(FMR, MVSC, SFMC, RMKM) encounter the ’Out-of-Memory’ problem. As a result, the time bars of these methods are omitted in the comparison graph.
@@ 4.2.3 Parameter Analysis
4.3 Extensional Study
We will further investigate whether the number of categories of the datasets can positively affect the accuracy of clustering by providing guidance for the assignation of the number of anchors; and whether it can be beneficial to the heuristic setting of the dimensions of each layer.

### v2

@@ 3.2.1 Optimization Algorithm
The decision variables for the objective function are $\{\mathbf{W}^{(v)}_{i}\}_{i=1}^{\delta},\mathbf{A}$, $\mathbf{Z}$ and $\boldsymbol{\alpha}$. As a result, we introduce a four-step iterative optimization method to approximate the optimal target mentioned in Eq. (6). In this approach, we keep three variables fixed while optimizing the remaining ones. The specific procedure is derived from the following detailed steps. Initialization. Our algorithm adopts a straightforward initialization by setting $\mathbf{Z}$ as an $m\times n$ matrix with the first $m\times m$ block deposited as identity matrix and the rest elements deposited to be 00. The anchor matrix $\mathbf{A}$ is initialized as identity matrix w.r.t. its dimension. And $\boldsymbol{\alpha}$ is initialized using mean values factorized by view count. The other decision variable $\{\mathbf{W}^{(v)}_{i}\}_{i=1}^{\delta}$ is neglected for initializing as it is the first to optimize. Update $\mathbf{W}^{(v)}_{o}$ with $\{\mathbf{W}^{(v)}_{i}\}_{i=1,i\neq o}^{\delta}$, $\mathbf{A}$, $\mathbf{Z}$ and $\boldsymbol{\alpha}$ fixed. Given $\mathbf{A}$, $\mathbf{Z}$, $\boldsymbol{\alpha}$ and the rest $\delta-1$ numbers of $\{\mathbf{W}^{(v)}_{i}\}_{i=1,i\neq o}^{\delta}$, the optimization problem in Eq. (6) with respect to $\mathbf{W}^{(v)}_{o}$ becomes:
@@
Denote the SVD of matrix $\mathbf{M}^{(v)}$ as $\mathbf{U}_{\mathbf{M}}\mathbf{D}_{\mathbf{M}}\mathbf{V}_{\mathbf{M}}^{\top}$ and the optimal solution of Eq. (9) is demonstrated as ${\mathbf{W}^{(v)}_{o}}^{\star}$. In this case, $\mathbf{D}_{\mathbf{M}}$ is a nonnegative diagonal matrix with each entry representing the singular value of $\mathbf{M}^{(v)}$, both $\mathbf{U}_{\mathbf{M}}$ and $\mathbf{V}_{\mathbf{M}}$ are unitary matrix with orthogonal entries. According to deduction in Ou et al. (2020), a closed-form resolution can be found by calculating ${\mathbf{W}^{(v)}_{o}}^{\star}=\mathbf{U}_{\mathbf{M}}sign(\mathbf{D}_{\mathbf{M}})\mathbf{V}_{\mathbf{M}}^{\top}$, where $sign(\cdot)$ is a function imposed on matrix with each element, which returns $1$ if the element is larger than 00. And it is obvious that in our case, $sign(\mathbf{D}_{\mathbf{M}})$ is an identity matrix. Accordingly, the optimal solution can be calculated as:
${\mathbf{W}^{(v)}_{o}}^{\star}=\mathbf{U}_{\mathbf{M}}\mathbf{V}_{\mathbf{M}}^{\top}$ (10)
By following the updating expression in Eq. (10), our algorithm updates $\mathbf{W}^{(v)}_{i},~{}i=1,2,\cdots,\delta$ in succession. Update $\mathbf{A}$ with $\{\mathbf{W}^{(v)}_{i}\}_{i=1}^{\delta}$, $\mathbf{Z}$ and $\boldsymbol{\alpha}$ fixed. When $\boldsymbol{\alpha}$ and $\{\mathbf{W}^{(v)}_{i}\}_{i=1}^{\delta}$ of each view are all prescribed, updating $\mathbf{A}$ with fixed $\mathbf{Z}$, which relocates the unified anchor matrix, can be termed as:
@@
where $\Phi=\sum_{v=1}^{p}\alpha_{v}^{2}{\mathbf{W}^{(v)}_{\delta}}^{\top}{\mathbf{W}^{(v)}_{\delta-1}}^{\top}\cdots{\mathbf{W}^{(v)}_{1}}^{\top}\mathbf{X}^{(v)}\mathbf{Z}^{\top}$ and $\Phi\in\mathbb{R}^{k\times m}$. Corresponding to the updating rule in Ou et al. (2020), it is straightforward to obtain the optimal solution for $\mathbf{A}$ as in Eq. (9). The multiplication of the left and right singular matrices of $\Phi$ costs $\mathcal{O}(km^{2})$. Update $\mathbf{Z}$ with $\mathbf{A}$, $\{\mathbf{W}^{(v)}_{i}\}_{i=1}^{\delta}$ and $\boldsymbol{\alpha}$ fixed. Given $\{\mathbf{W}^{(v)}_{i}\}_{i=1}^{\delta}$, the anchors are separately lodged in $p$ different anchor spaces of the same dimension $k$ and are reunited by the optimized anchor matrix $\mathbf{A}$, which is shared amongst views. This subtask of updating $\mathbf{Z}$ is equivalent to the process of constructing the bipartite graph from the anchor space to the original feature space. The optimization problem in Eq. (6) concerning $\mathbf{Z}$ can be written as:
@@
for $j=1,2,\cdots,n$, where the parameter for quadratic term is $\mathbf{Q}=2\sum_{v=1}^{p}{\alpha_{v}^{2}}\mathbf{I}$ and the first order polynomial term ${\boldsymbol{q}^{\top}}=-2\sum_{v=1}^{p}{\alpha_{v}^{2}}{\mathbf{X}^{(v)}_{:,j}}^{\top}\mathbf{P}$. These are several QP(quadratic programming) problems with domain constraints. As each column of $\mathbf{Z}$ is an $m$ dimensional vector, and there are $n$ sub-problems identical to each other, the overall time cost of optimising $\mathbf{Z}$ is $\mathcal{O}(nm^{3})$. Update $\boldsymbol{\alpha}$ with $\mathbf{Z}$, $\mathbf{A}$ and $\{\mathbf{W}^{(v)}_{i}\}_{i=1}^{\delta}$ fixed. When variables $\mathbf{Z}$, $\mathbf{A}$ and $\{\mathbf{W}^{(v)}_{i}\}_{i=1}^{\delta}$ are fixed, the optimization problem in Eq. (6) with regard to the variable $\boldsymbol{\alpha}$ automatically learns the coefficients of different views, the procedure is assembled below:
@@ 4.1.1 A Brief Introduction of Datasets
For in-depth experimental analysis, we conduct a series of tests over 10 real-world datasets, varying in multiple domains. The datasets used in our experiments are WebKB  [^1] http://lig-membres.imag.fr/grimal/data.html, BDGP  [^2] https://www.fruitfly.org/, MSRCV1 Winn and Jojic (2005), Caltech7-2view  [^3] http://www.vision.caltech.edu/Image_Datasets/Caltech101/, NUS-WIDE-10  [^4] https://lms.comp.nus.edu.sg/wp-content/uploads/2019/research/nuswide/NUS-WIDE.html, Caltech101-7 ${}^{3}$, Caltech101-20 ${}^{3}$, NUS-WIDE ${}^{4}$, SUNRGBD  [^5] https://rgbd.cs.princeton.edu/, YoutubeFace  [^6] https://www.cs.tau.ac.il/ wolf/ytfaces/. These datasets vary in broad ranges of application background. In specific, WebKB is a dataset that includes web pages from computer science departments of various universities and they are categorized into 5 categories (Student, Faculty, Department, Course, Project). The SUNRGBD dataset contains 10335 real RGB-D images of room scenes. The MSRCV1 dataset from Microsoft Research in Cambridge contains 7 classes of images with coarse pixel-wise labeling. In Table 2, we summarize the detailed information of these benchmark datasets.
@@ 4.1.2 Experimental Settings
We use the initialization specified in Algorithm 1. In practice, we simply adopt the number of clusters $k$ as the value of the dimension of embedding space and the value of anchor count. For each view, the hierarchical feature descent rate is different according to their original dimensions. However, it is rather practicable as it performs well by directly slicing the dimension to equal sections: $l_{o}-l_{o-1}=l_{o+1}-l_{o},\forall o\in[1,\delta]$. The best depths among $[1,2,\cdots,5]$ of the datasets WebKB, BDGP, MSRCV1, Caltech7-2view, NUS-WIDE-10, Caltech101-7, Caltech101-20, NUS-WIDE ${}^{4}$, SUNRGBD, YoutubeFace are 5, 5, 5, 5, 4, 4, 5, 5, 2, respectively. To be specific, the experiments are conducted in MATLAB environment on PC with Intel(R) Core(TM) i7-6800K 3.40GHz CPU and 128GB RAM.
@@ 4.1.3 Compared Algorithms
RMKM LMVSC AMGL BMVC FMR MVSC SFMC FPMVS-CAG EMKMC proposed
#Parameter 1 2 0 3 2 3 1 0 $p$ 1
As shown in Table.3, most of algorithm have more than one hyper-parameters. It is encouraged to use learnable parameters which can adaptively collaborate with the optimization of other decision variables to boost the performance whereas spare the acquisition of prior knowledge. In our method, the learnable vector of parameters $\boldsymbol{\alpha}=[\alpha_{1},\alpha_{2},\cdots,\alpha_{p}]^{\top}$ represents the weights adherent to different views, and they are optimized in each iteration. However, for the compared algorithm EMKMC, each view has a hyper-parameter, so the number of parameters is equal to the number of views, which we denote by $p$. As a result, the searching space for best parameter is $\epsilon^{p}$ where $\epsilon$ denotes the size of the candidate set of the hyper-parameter. These unlearnable parameters dramatically increases computation time of EMKMC.
@@ 4.2 Comparison With State-of-the-Art Algorithms
In this section, we conduct a comparative analysis between our algorithm and several state-of-the-art MVC algorithms to validate its effectiveness. We evaluate the performance of our algorithm using 9 widely-used methods, considering accuracy (ACC), Normalized Mutual Information (NMI), and purity as evaluation metrics. The results are presented in Table 4. It shows that our findings entail most of the best practices and recommendations that can improve multi-view learning performance. To illustrate the effectiveness of the anchors, we conduct t-Distributed Stochastic Neighbor Embedding (t-SNE) experiment and mark the anchors with red crosses. Due to space constraints, we show the clustering effect for WebKB, MSRCV1 and Caltech7-2views in Figure 3. We observe that a red cross is located in each cluster, which empirically verifies the effectiveness of the optimization of anchors in the objective function.
(a) WebKB
(b) MSRCV1
(c) Caltech7-2views
Figure 3: The t-SNE plot of our proposed method on 3 benchmark datasets
@@ 4.2.1 Running Time Comparison
The running time of compared algorithms on benchmark datasets are shown in Figure 4. Due to the excessive time consumed by some of the algorithms, we use the logarithms of the time as the vertical axis of the coordinate.
@@
It is worth noticing that for large-scale datasets, namely NUS-WIDE, SUNRGBD and YoutubeFace, some algorithms(FMR, MVSC, SFMC, RMKM) encounter the ’Out-of-Memory’ problem. As a result, the time bars of these methods are omitted in the comparison graph. We observe that the time cost of EMKMC is comparable to ours, but the total runtime cost is still high owing to the aforementioned search space for the best hyper-parameters.
@@ 4.2.3 Parameter Analysis
4.3 Conclusion
Our work focuses on multi-view clustering and explore ways to increase the effectiveness of anchor-based subspace clustering by tackling the discrepancy and deploying the dependency of different views. We will further investigate whether the number of categories of the datasets can positively affect the accuracy of clustering by providing guidance for the assignation of the number of anchors; and whether it can be beneficial to the heuristic setting of the dimensions of each layer.